In a company, out of the 7 applicants for 3 positions, 3 are males and 4 are females. In how many ways will 1 males and 2 female be chosen? Thank you.
Answer by oscargut(2103)

Answer: 3* (4C2) = 3*6 = 18

I can help you with all of your questions at: mthman@gmail.com
Thanks